OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

it's not a system

OpenStudy (anonymous):

y=2x-6....................1 4x-2y=12.................2 ....................................... sub 1 in 2 4x-2(2x-6)=12 4x-4x+12=12 Its clear there is no such Q... regard :P

OpenStudy (anonymous):

excuse me mmathhenius?!?! watch your self i did nothing wrong

OpenStudy (ash2326):

Mmathgenius be nice, help her out.

OpenStudy (anonymous):

I thought mathgenius was calling her retarded for not knowing how to do this, but he is saying it was just explained to her I guess.

OpenStudy (anonymous):

This is not my work it is my sisters and i do not have any idea how to solve this u dont have to answer to this questiopn u can go to someone eleses ok i was not begging u to answer this

OpenStudy (anonymous):

This is a inconsistent system, two parallel straight lines.

OpenStudy (ash2326):

coral1998 leave it, we'll help you:)

OpenStudy (anonymous):

It is fine i will not ask any more idk why some ppl have to be so rude tho :\ thANK ALL OF U WHO HELPED ME! that will be all.....

OpenStudy (farmdawgnation):

coral1998, sorry about that. I've issued that user an official warning. If they do something like that again - file an Abuse Report and we'll handle it.

OpenStudy (anonymous):

Ok thank u :) and u use big word lol :p jkjk